# TurBLiMP: A Turkish Benchmark of Linguistic Minimal Pairs

### v1

@@ 7 Results
We see that the monolingual models BERTurk and cosmosGPT tend to outperform their multilingual counterparts. Their performance is comparable to the best multilingual models EuroLLM and Gemma 3 12B. BERTurk is the only model that shows a strong cross-phenomenon correlation with human acceptability ratings, as illustrated in Figure 2. This is worth noting given that BERTurk is the only masked language model that we have tested. None of the other models had a statistically significant correlation in either direction.
@@
Finally, the Goldfish model series reveals the effect of training data size on performance. Models with larger training data sizes typically achieve better performance, though some counter-intuitive patterns emerge near random-chance levels. While more data generally improves learning, this pattern does not hold when acceptable sentences are consistently shorter than unacceptable ones.
@@ 7.1 Effect of Word Order
Although SOV is the canonical word order in Turkish, Slobin and Bever (1982) found that 52% of utterances in their spontaneous adult speech corpus deviate from this order. Similarly, Türk et al. (2022) reported that only 59.5% of sentences in the BOUN Universal Dependencies Treebank follow SOV. Notably, they identified two different word orders as the second most frequent, highlighting how Turkish word order patterns can vary largely between spoken and written language. Both studies, however, agree that VOS is the least attested.
Native-speaker acceptability judgments reflect that SOV had the highest mean rating difference for both transitive and ditransitive sentences, in line with spoken and written corpus frequencies. The second-highest mean acceptability rating difference for transitive paradigms was the SVO word order, while it was OVS for the ditransitive ones. These are also the second-most-frequent word orders reported by Slobin and Bever (1982) and Türk et al. (2022) respectively. In transitive sentence ratings, VOS is not found to be the most challenging word order. This suggests that a rare word order does not inherently hinder people’s ability to identify acceptable sentences. Speakers seem to tolerate non-canonical word orders more readily in transitives than in ditransitives. One interpretation may be that case differences are easier to spot in transitive sentences due to fewer arguments.
Data Metric SOV SVO OSV OVS VSO VOS Human Slobin and Bever Freq. 48.0 25.0 8.0 13.0 6.0 0.0 BOUN Treebank Freq. 59.5 5.9 4.5 22.4 6.8 0.9 Arg. Str. Tran. $\Delta$ 1.98 1.92 1.81 1.70 1.68 1.75 Arg. Str. Ditr. $\Delta$ 1.81 1.56 1.47 1.62 1.45 1.29 EuroLLM Arg. Str. Tran. $\Delta$ 5.24 3.40 2.77 4.39 2.42 2.83 Arg. Str. Ditr. $\Delta$ 7.83 6.13 5.66 7.11 5.30 4.04 BERTurk Arg. Str. Tran. $\Delta$ 13.41 10.24 11.92 13.44 7.87 10.16 Arg. Str. Ditr. $\Delta$ 15.33 8.94 11.21 14.18 7.86 4.74
Table 3: Word order performance comparison between human judgments and best models. The white-blue gradient represents mean acceptability differences (low to high) for each row, while the white-yellow gradient reflects corpus frequency.
We see the opposite trend for model evaluations with EuroLLM being particularly sensitive to non-canonical word orders in transitive sentences. BERTurk remains robust to all word orders, showing only a pronounced drop for the rare VOS paradigm in the ditransitive condition. For both transitive and ditransitive sentences, models show high mean log probability differences on OVS word orders. This suggests that model performances align more closely with word order statistics from the BOUN treebank than with those from the spoken language corpus by Slobin and Bever (1982).
@@ 7.2 Effect of Subordination
Table 4 displays human and model performance on four subordination paradigms compared to a non-subordinated baseline. In Turkish, subordinate clauses can be finite or non-finite. However, finite subordinate clauses are much less frequent than non-finite ones (Göksel and Kerslake, 2005). For non-finite subordination, we consider three different subordinating suffixes: -DIK, -(y)IncA, and -(y)ken. -DIK forms nominal subordinate clauses while the latter two form adverbial ones.
@@ 8 Conclusion
TurBLiMP [^1] https://github.com/ezgibasar/turblimp provides the first comprehensive evaluation of language models’ syntactic capabilities for Turkish. We find that larger model sizes generally correlate with higher accuracy, with some exceptions. Considerably smaller monolingual language models often outperform their larger multilingual counterparts and perform on par with the best multilingual models. This finding corroborates patterns attested in other syntactic benchmarks Taktasheva et al. (2024); Jumelet et al. (2025). The strong performance of monolingual models highlights the importance of language-specific training for reliable models. Cases where smaller models outperformed larger ones also suggest that scaling alone cannot explain model behavior as far as linguistic evaluations are concerned.
@@
While TurBLiMP offers a comprehensive evaluation of key linguistic phenomena in Turkish, there are also several limitations to acknowledge. In this paper, we evaluated minimal pair acceptability using sequence log probabilities for each sentence. However, our approach represents only one of several valid methods for assessing language models on acceptability benchmarks. For example, Song et al. (2025) show that prompting for ‘meta-linguistic’ grammaticality judgments can result in better performance than comparing string probabilities directly. However, they do show that this ‘introspective’ approach has its limitations, and the optimal way of evaluating linguistic ability in LMs remains an open debate (Hu et al., 2024).
Warstadt et al. (2020) define paradigms as minimal pair types and phenomena as broader linguistic categories. Unlike other BLiMP benchmarks that include multiple paradigms under each phenomenon, TurBLiMP currently includes only one paradigm per phenomenon (with the exception of Argument Structure). Splitting some phenomena into multiple paradigms could enable more granular assessment. For a broader coverage of linguistic structures in Turkish, future work could also incorporate additional phenomena into the benchmark.
@@
Some of the models we tested had both base and instruction-tuned variants. We excluded instruction-tuned variants as the English-based tuning was unlikely to improve performance on Turkish linguistic tasks. However, this design choice means we cannot speak to potential transfer effects from instruction tuning.
Our word order experiments focused exclusively on sentences with explicit subjects, omitting pro-drop constructions. Given Türk et al. (2022)’s finding that subjectless sentences exceed canonical SOV order frequency in the BOUN treebank, future work should investigate these prevalent but untested configurations.
@@ Appendix B Acceptability Judgment Collection
Figure 3 provides a screenshot of the survey carried out on the Qualtrics platform. All participants gave their informed consent before starting the survey and agreed that their anonymous responses can be made publicly available. Students received extra credit in exchange for their participation.

### v2

@@ 7 Results
We see that the monolingual models BERTurk and cosmosGPT tend to outperform their multilingual counterparts. Their performance is comparable to the best multilingual models EuroLLM and Gemma 3 12B. BERTurk is the only model that shows a strong cross-phenomenon correlation with human acceptability ratings, as illustrated in Figure 2. This is worth noting given that BERTurk is the only masked language model that we have tested. None of the other models had a statistically significant correlation in either direction. While the reported correlations are based on sequence log probabilities, we also experimented with using Syntactic Log-Odds Ratio (SLOR) (pauls-klein-2012-large) scores to rule out the possibility that poor correlations were an artifact of the chosen metric. As an acceptability measure, SLOR normalizes sequence probabilities by controlling for sentence length and word frequency. SLOR-based correlations are provided in Appendix D.
@@
Finally, the Goldfish model series reveals the effect of training data size on performance. Models with larger training data sizes typically achieve better performance, though some counter-intuitive patterns emerge near random-chance levels. While more data generally improves learning, this pattern does not hold when acceptable sentences are consistently shorter than unacceptable ones. In addition to the reported model results, a supplementary analysis of surface-level confounds such as sentence length and subword counts can be found in Appendix E.
@@ 7.1 Effect of Word Order
Data Metric SOV SVO OSV OVS VSO VOS Human SLOBIN1982229 Freq. 48.0 25.0 8.0 13.0 6.0 0.0 BOUN Treebank Freq. 59.5 5.9 4.5 22.4 6.8 0.9 Arg. Str. Tran. $\Delta$ 1.98 1.92 1.81 1.70 1.68 1.75 Arg. Str. Ditr. $\Delta$ 1.81 1.56 1.47 1.62 1.45 1.29 EuroLLM Arg. Str. Tran. $\Delta$ 5.24 3.40 2.77 4.39 2.42 2.83 Arg. Str. Ditr. $\Delta$ 7.83 6.13 5.66 7.11 5.30 4.04 BERTurk Arg. Str. Tran. $\Delta$ 13.41 10.24 11.92 13.44 7.87 10.16 Arg. Str. Ditr. $\Delta$ 15.33 8.94 11.21 14.18 7.86 4.74
Table 3: Word order performance comparison between human judgments and best models. The white-blue gradient represents mean acceptability differences (low to high) for each row, while the white-yellow gradient reflects corpus frequency. Frequency values represent the percentage of different word orders within each corpus. For human judgments, mean differences are calculated using z-score–transformed acceptability ratings, whereas for model evaluation, the mean difference reflects the difference in sequence log probabilities.
Although SOV is the canonical word order in Turkish, SLOBIN1982229 found that 52% of utterances in their spontaneous adult speech corpus deviate from this order. Similarly, 10.1007/s10579-021-09558-0 reported that only 59.5% of sentences in the BOUN Universal Dependencies Treebank follow SOV. Notably, they identified two different word orders as the second most frequent, highlighting how Turkish word order patterns can vary largely between spoken and written language. Both studies, however, agree that VOS is the least attested.
Native-speaker acceptability judgments reflect that SOV had the highest mean rating difference for both transitive and ditransitive sentences, in line with spoken and written corpus frequencies. The second-highest mean acceptability rating difference for transitive paradigms was the SVO word order, while it was OVS for the ditransitive ones. These are also the second-most-frequent word orders reported by SLOBIN1982229 and 10.1007/s10579-021-09558-0 respectively. In transitive sentence ratings, VOS is not found to be the most challenging word order, contrary to what might be expected based on attested corpus statistics. This suggests that a rare word order does not inherently hinder people’s ability to identify acceptable sentences. Speakers seem to tolerate non-canonical word orders more readily in transitives than in ditransitives. One interpretation may be that case differences are easier to spot in transitive sentences due to fewer arguments.
We see the opposite trend for model evaluations with EuroLLM being particularly sensitive to non-canonical word orders in transitive sentences. BERTurk remains robust to all word orders, showing only a pronounced drop for the rare VOS paradigm in the ditransitive condition. For both transitive and ditransitive sentences, models show high mean log probability differences on OVS word orders. This suggests that model performances align more closely with word order statistics from the BOUN treebank than with those from the spoken language corpus by SLOBIN1982229.
@@ 7.2 Effect of Subordination
Table 4 displays human and model performance on four subordination paradigms compared to a non-subordinated baseline. In Turkish, subordinate clauses can be finite or non-finite. However, finite subordinate clauses are much less frequent than non-finite ones (göksel2005turkish). For non-finite subordination, we consider three different subordinating suffixes: -DIK, -(y)IncA, and -(y)ken. -DIK forms nominal subordinate clauses while the latter two form adverbial ones.
@@ 8 Conclusion
TurBLiMP [^1] https://github.com/ezgibasar/turblimp provides the first comprehensive evaluation of language models’ syntactic capabilities for Turkish. We find that larger model sizes generally correlate with higher accuracy, with some exceptions. Considerably smaller monolingual language models often outperform their larger multilingual counterparts and perform on par with the best multilingual models. This finding corroborates patterns attested in other syntactic benchmarks taktasheva-etal-2024-rublimp; jumelet2025multiblimp10massivelymultilingual. The strong performance of monolingual models highlights the importance of language-specific training for reliable models. Cases where smaller models outperformed larger ones also suggest that scaling alone cannot explain model behavior as far as linguistic evaluations are concerned.
@@
While TurBLiMP offers a comprehensive evaluation of key linguistic phenomena in Turkish, there are also several limitations to acknowledge. In this paper, we evaluated minimal pair acceptability using sequence log probabilities for each sentence. However, our approach represents only one of several valid methods for assessing language models on acceptability benchmarks. For example, song2025languagemodelsfailintrospect show that prompting for ‘meta-linguistic’ grammaticality judgments can result in better performance than comparing string probabilities directly. However, they do show that this ‘introspective’ approach has its limitations, and the optimal way of evaluating linguistic ability in LMs remains an open debate (doi:10.1073/pnas.2400917121).
warstadt-etal-2020-blimp-benchmark define paradigms as minimal pair types and phenomena as broader linguistic categories. Unlike other BLiMP benchmarks that include multiple paradigms under each phenomenon, TurBLiMP currently includes only one paradigm per phenomenon (with the exception of Argument Structure). Splitting some phenomena into multiple paradigms could enable more granular assessment. For a broader coverage of linguistic structures in Turkish, future work could also incorporate additional phenomena into the benchmark.
@@
Some of the models we tested had both base and instruction-tuned variants. We opted to exclude instruction-tuned versions based on our assumption that English-based tuning was unlikely to improve performance on Turkish grammatical evaluation tasks. Prior work chirkova-nikoulina-2024-zero-shot has shown that instruction tuning in English can degrade fluency in non-English languages, and we also observed performance losses when experimenting with the instruction-tuned counterparts of several models. For instance, the average accuracy of Gemma 3 (4B) Instruct was 4.4% lower than its base variant. That said, this experimental choice limits our ability to comment on potential transfer effects from instruction tuning.
Our word order experiments focused exclusively on sentences with explicit subjects, omitting pro-drop constructions. Given 10.1007/s10579-021-09558-0’s finding that subjectless sentences exceed the canonical SOV order frequency in the BOUN treebank, future work should investigate these prevalent but untested configurations.
Acknowledgments
We are thankful to the anonymous reviewers for their helpful comments and to Ümit Atlamaz for his assistance in recruiting participants. We also thank the survey participants for their time and insight. Ezgi Başar received funding from the Erasmus Mundus Masters Programme in Language and Communication Technologies, EU grant no. 2019-1508. Francesca Padovani, Jaap Jumelet and Arianna Bisazza were supported by the Talent Programme of the Dutch Research Council (grant VI.Vidi.221C.009).
@@ Appendix B Acceptability Judgment Collection
Figure 3 provides a screenshot of the survey carried out on the Qualtrics platform. All participants gave their informed consent before starting the survey and agreed that their anonymous responses can be made publicly available. Students received extra credit in exchange for their participation. Participation was on a voluntary basis for both students and non-students, and they were explicitly informed that no financial compensation would be provided. The survey was distributed within our close networks.
@@ Appendix C Acceptability Judgment Statistics
Appendix D SLOR-based Human Correlations
Model Pearson’s $r$ $p$-value
Goldfish 1000MB 0.384 0.1422
BERTurk 0.557 0.0251
CosmosGPT 0.414 0.1105
Gemma 3 (4B) 0.400 0.1247
Qwen 2.5 0.233 0.3845
Llama 3.1 0.283 0.2881
Aya Expanse 0.454 0.0771
Gemma 2 0.358 0.1730
EuroLLM 0.357 0.1752
Gemma 3 (12B) 0.385 0.1411
Table 10: SLOR-based correlation coefficients for various models and their corresponding $p$-values. The statistically significant result ($p<0.05$) is indicated in boldface.
Appendix E Phenomenon-Agnostic Factors
Figure 4: $\beta$ coefficients fitted for the BERTurk, EuroLLM, Goldfish, and Qwen 2.5 models with sentence length and subword count differences as the predictors.
To understand general factors influencing model performance beyond phenomena-specific evaluations, we analyzed two potential predictors across all minimal pairs: (1) the difference in sentence lengths between acceptable and unacceptable sentences, and (2) the difference in subword counts for acceptable and unacceptable cue words.
Subword counts refer to the number of smaller units, or subwords, into which a word is segmented by a tokenizer. Depending on the tokenizer design, the resulting subwords may approximate word or morpheme boundaries, but they typically do not perfectly align with either. The subword count difference metric captures the disparity in tokenization granularity between the acceptable and unacceptable cue words and shows whether the acceptable word is split into fewer or more subword units than the unacceptable word.
Using these two predictors, we conduct a linear modeling experiment to predict log probability differences obtained by four different models. We include the two best-performing models (EuroLLM and BERTurk) as well as the two worst-performing models (Goldfish and Qwen 2.5). To allow for better comparability, we opt for the largest Turkish Goldfish model trained on 1000MB of data.
Figure 4 displays the standardized $\beta$ coefficients for both sentence length differences and subword count differences across all four models. We also include each model’s corresponding $R^{2}$ value, indicating how well these predictors explain the variance in log probability differences.
$R^{2}$ values are remarkably low for all the models, which means that these two factors alone are not adequate predictors of model behavior. For the purposes of our benchmark, this is the desired outcome as we want model log probabilities to reflect grammatical judgments rather than being confounded by these surface-level features.
Sentence length
Autoregressive and masked models show sentence length effects in opposite directions. In the case of autoregressive models, the model is more likely to incorrectly prefer the unacceptable sentence if the acceptable sentence is longer than its unacceptable counterpart. This indicates that autoregressive models are biased to prefer shorter sequences. BERTurk, which is the only masked language model in our model pool, shows the opposite pattern. As the acceptable sentence gets longer than the unacceptable one, BERTurk is more likely to make the correct prediction. This is possibly due to longer sentences providing more contextual clues for masked token prediction.
Reflecting back on the previously reported accuracy results, we can note that acceptable sentences are typically shorter than unacceptable ones in NPI Licensing, Passives and Quantifiers while the opposite is true in Determiners. Autoregressive models perform better on NPI Licensing, Passives and Quantifiers while BERTurk exhibits slightly better performance for Determiners. This behavioral pattern aligns with the results of our linear model.
Subword counts
Based on prior work (goldman-etal-2024-unpacking; jumelet2025multiblimp10massivelymultilingual), we expect model performance to degrade with increasing subword count differences. The Goldfish model shows a negative subword count coefficient, behaving in line with our hypothesis. A negative coefficient indicates that the model is more likely to make the wrong prediction if the acceptable cue word is split into more subwords than the unacceptable one. However, contrary to our expectations, all other models show a weak but consistent positive effect. To help us make sense of this counterintuitive finding, we conduct a follow-up analysis examining how each model’s tokenizer aligns with morphological boundaries in our stimuli.
Utilizing the morphological inflection pipeline by akin2007zemberek, we construct a small dataset of 400 morphologically segmented Turkish words. The dataset is organized into four categories, each targeting morphemes that frequently appear as part of the cue words in our benchmark.
  • (Verb + Nominalizer + Possessive + Case): Contains 100 verb forms with the nominalizer -DIK followed by possessive markers and case endings (e.g., unuttuğunu $\rightarrow$ [unut, tuğ, un, u]). This category evaluates the segmentation of nominalized verbs.
  • (Verb + TAM + Person): Includes 100 finite verbs segmented into stems, tense/aspect/modality (TAM) markers, and person markers (e.g., olacaktılar $\rightarrow$ [ol, acak, , lar]), targeting inflectional morphology.
  • (Noun + (Plural) + Possessive + Case): Comprises 100 possessed nouns decomposed into stems, optional plural markers (-lAr), possessive markers, and case endings (e.g., elmalarıma $\rightarrow$ [elma, lar, ım, a]).
  • (Noun + (Plural) + Case): Features 100 simpler noun forms with optional plural and case markers (e.g., elmalara $\rightarrow$ [elma, lar, a]), providing a baseline for bare nominal inflection.
arnett-bergen-2025-language previously investigated the morphological alignment of tokenizers across typologically diverse languages to investigate whether a lack of alignment could explain performance gaps in language models. They found no evidence to suggest that morphological alignment plays a significant role. However, their operationalization of morphological alignment differs from ours as far as Turkish is concerned. While they assess whether the tokenizer separates the stem from all other suffixes combined, we investigate whether the tokenizer identifies all the different morpheme boundaries.
We evaluate the models based on their ability to approximate gold-standard morpheme boundaries for the 400 words in our dataset. To that end, we use four different metrics to quantify morphological alignment. These four metrics are: (1) the average Damerau-Levenshtein distance (damerau) between the gold morphemes and the subwords obtained by the model, (2) the proportion of undersegmented words which refers to cases where the model produced fewer segments than the gold data, (3) the proportion of oversegmented words, and (4) the proportion of times when the model produced an output identical to the gold data.
Tokenizer Avg. Undersegm. Oversegm. Exact
Distance Items Items Matches
BERTurk 2.48 82.2 0.8 5.8
EuroLLM 2.36 24.8 27.5 2.5
Goldfish 1.62 83.0 3.5 9.8
Qwen 2.5 5.50 7.0 54.2 0.5
Table 11: Morphological alignment results.
Table 11 illustrates each tokenizer’s alignment to gold-standard morpheme boundaries. We observe no correlation between subword count coefficients and the latter three metrics (undersegmentation rate, oversegmentation rate, and exact match proportion). Interestingly, the ranking of models by Damerau-Levenshtein distance perfectly mirrors their ranking by subword count coefficients.
The two best-performing models and the two worst-performing models do not cluster together based on either their subword count coefficients or morphological alignment scores. This reinforces the observation that performance differences cannot be attributed to tokenizer behavior alone.